OpenStudy (valiyuh):

a ramp is 3 m above the ground and is 4.5 m long. john uses force of 500 n to lift the box from the ground straight up to the top of the ramp, while mike uses a force of 250 n t push the same box up the complete length of the ramp. what is the percent efficiency of the ramp?

OpenStudy (anonymous):

the work done by John= 500* 3=1500 J the work done by Mike=250*sqrt(4.5^2+3^2)=250*5.53=1382.5 J %=(1500-1382.5)/1500*100=7.8

OpenStudy (anonymous):